### v1

# MistralBSM: Leveraging Mistral-7B for Vehicular Networks Misbehavior Detection

Abstract
Vehicular networks are exposed to various threats resulting from malicious attacks. These threats compromise the security and reliability of communications among road users, thereby jeopardizing road and traffic safety. One of the main vectors of these attacks within vehicular networks is misbehaving vehicles. To address this challenge, we propose deploying a pretrained Large Language Model (LLM)-empowered Misbehavior Detection System (MDS) within an edge-cloud detection framework. Specifically, we fine-tune Mistral-7B, a state-of-the-art LLM, as the edge component to enable real-time detection, whereas a larger LLM deployed in the cloud can conduct a more comprehensive analysis. Our experiments conducted on the extended VeReMi dataset demonstrate Mistral-7B’s superior performance, achieving 98% accuracy compared to other LLMs such as LLAMA2-7B and RoBERTa. Additionally, we investigate the impact of window size on computational costs to optimize deployment efficiency. Leveraging LLMs in MDS shows interesting results in improving the detection of vehicle misbehavior, consequently strengthening vehicular network security to ensure the safety of road users.
Index Terms:
Vehicular network, Large Language Models (LLMs), misbehavior detection, intrusion detection
I Introduction
The evolution of vehicular networks [1], commonly known as vehicle-to-everything (V2X), has been marked by continuous developments in communication and transportation technologies. Originally conceived as a means of improving vehicle safety through the exchange of information between vehicles and the infrastructure, research expanded into areas such as cooperative driving, advanced scenarios to support automated driving, and the integration of vehicle networks with emerging technologies such as 5G and beyond [2] and the Internet of Things (IoT) [3]. Today, as the potential of connected autonomous vehicles promises to redefine the future of private and public transportation, the evolution of vehicular networks research continues apace, particularly with regard to cybersecurity issues[4]
Indeed, vehicular networks are exposed to various threats resulting from malicious attacks [5]. These threats compromise the security and reliability of communications among road users, thereby jeopardizing road and traffic safety. One of the main vectors of these attacks within vehicular networks is misbehaving vehicles [6]. Misbehaving vehicles can launch denial-of-service (DoS) attacks which flood the network with malicious traffic, disrupting communication between vehicles [7]. As a result, this disruption can lead to communication breakdowns, making vehicles unaware of each other and disrupting the operation of both traffic systems and vehicular safety applications. In addition, message tampering and integrity attacks can mislead on-board control systems, compromising vehicle safety. In this scenario, misbehaving vehicles deliberately transmit inaccurate or erroneous messages despite normal hardware and software functioning [8]. Vehicle misbehavior can result in significant repercussions, including serious traffic disruptions or even accidents. Hence, detecting misbehavior is crucial in vehicular communication systems. However, it is a complex challenge due to the dynamic nature of traffic environments and the diverse forms misbehavior can take. Vehicles interact unpredictably, leading to various behaviors that may differ from expected patterns. Additionally, the large volume of data generated by vehicles adds complexity. Analyzing this data in real-time requires advanced algorithms and computing resources.
Machine learning (ML) has recently emerged as an interesting technology in the development of misbehavior detection systems (MDS). ML enables MDS to adapt and learn from historical data, continuously improving its ability to identify abnormal or malicious behavior. [9]. With the recent advancements in large language models (LLMs) there is an exciting opportunity to further bolster the capabilities of MDS. Indeed, models, such as the GPTs [10], LLAMA-2[11] and BERT[12] have demonstrated remarkable capabilities in various natural language processing tasks, but also showed interesting capacities in network security tasks [13], including for vehicular networks security [14, 15].
The potential of LLMs in vehicle misbehavior detection stems from the fact that LLMs can facilitate the extraction of insights from textual data of vehicle communications. Through semantic analysis and contextual understanding, these models can identify linguistic features that can signify potential abnormal behavior patterns. However, using LLMs for vehicle misbehavior detection poses significant challenges, notably in privacy and resource constraints. Given their extensive parameter counts, often in the tens of billions or even trillions, LLMs are naturally suited for cloud deployment. Nevertheless, transferring all vehicle communication data to cloud infrastructure poses potential risks for real-time misbehavior detection and raises concerns about data privacy.
Recently, a new large-scale model named Mistal-7B [16] has been introduced, rivaling existing models with fewer parameters thanks to various architectural choices. This opens the possibility to use an LLM-empowered MDS architecture with edge-cloud collaboration, where a light LLM Model such as Mistral-7B resides at the edge (e.g., at a roadside unit), and a bigger LLM resides in the cloud. Edge LLMs can perform initial vehicle misbehavior detection in real-time, providing immediate feedback and alerts to vehicle systems. Cloud LLMs can complement this by conducting deeper analysis, incorporating additional contextual information from a broader, and possibly multi-modal dataset stored in the cloud (e.g., multimodal traffic information from the intelligent transportation system). Using the edge LLM has the advantage of processing vehicular communication data locally, minimizing the need to transmit raw data to the cloud and preserving user privacy. Cloud LLMs can be utilized for more general analysis that uses edge LLMs insights, and other data external to the communication system, ensuring compliance with privacy regulations and minimizing the risk of data privacy breaches.
In this article, we explore the use of Mistral-7B as an edge LLM to be located at a roadside unit, and we explore the capacity of the LLM to classify each vehicle as normal or misbehaving. The main contributions of this article are as follows:
  • We use Mistral-7B for misbehavior detection based on the analysis of observed basic safety massages (BSMs).
  • We transpose a vehicular misbehavior detection task into a sequence classification by converting windows of BSMs originating from a specific vehicle into text prompts to be classified by a language model.
  • We fine-tune Mistral-7B as a sequence classifier to assess a vehicle behavior nature through a sequence of its broadcasted BSMs, resulting in MistralBSM.
  • We assess Mistral-7B’s performance compared to other prominent open-source LLMs, namely LLAMA2-7B and BERT based architecture, in the context of MDS, and we demonstrate its superior results.
The rest of this paper is organized as follows. Section II introduces the related works of misbehavior detection methods in vehicular networks. Section III introduces a background about vehicular networks architecture, LLMs and low rank adaptation for LLM fine-tuning. In Section IV, we define our system model, the proposed solution and introduce Mistral for misbehavior detection. in Section V, we do a performance evaluation of our model and compare our results.
II Related works
II-A Vehicular misbehavior detection
Detecting misbehavior is widely recognized as a critical task in V2X security. ML-based approaches are commonly favored due to their distinctive ability to model and classify misbehavior, surpassing other rule-based strategies.
In [17], various models including Random Forest, Decision Tree, Ada Boost Classifier, Ridge Classifier, Naive Bayes, and XGBoost have been evaluated on the CAN-OITDS dataset for multi-class classification involving four distinct attack types. The results showed an advantage of XGBoost. The authors of [18], also employed ML technique for misbehavior detection. The proposed framework enhances the detection performance by including movement and position plausibilities with the BSM features in the input of the classifier. The considered algorithm includes SVM, KNN, Naive Bayes, Random Forest, ensemble boosting and voting ensemble. The primary drawback of such models is their limited ability to adequately handle the temporal aspect of specific attacks aimed at V2X communication.
To address this gap, some researchers considered using LSTM based architectures. For example, in [19], [20] and [21], LSTM and CNN were combined to detect misbehavior, yielding promising results. On the other hand, in [22] and [23], considered the misbehavior detection as time-series anomaly detection problem that could be modeled by a Markov Decision process where each decision is based on the previous ones. They proposed a DQN-based classifier where the agent q-network is formed by LSTM layers and a fully connected layer. Particularly, in [22], the proposed framework uses KNN as an unsupervised algorithm for labelling BSMs. Thereafter, the labeled data is used by the DQN agent to learn the optimal policy. The effectiveness of these techniques underscores the importance of considering the temporal dimension in detecting misbehavior within vehicular networks, where LSTM has demonstrated efficacy. However, interest has shifted with the introduction of transformers and LLMs, as they enable efficient processing of sequenced data.
II-B LLMs for network security
LLMs have showcased remarkable capabilities that extend beyond natural language processing (NLP), generating considerable interest across diverse domains. Notably, they have gained interest in network security, with encoder-based transformers being particularly renowned.
For example, BERT served as the foundation for log anomaly detection systems as described in ([24], [13], [25]). In [24], BERT was trained on normal log key sequences extracted from raw network logs to predict the next masked key. Subsequently, during the detection phase, the model predicted normal keys with high precision, while abnormal keys exhibited low detection accuracy. Conversely, in [13], the authors fine-tuned BERT using a fully connected classifier, where the input consists of a time window comprising a sequence of parsed log data. In [25], log anomaly detection is achieved by fine-tuning LongFormer, a transformer based on BERT, exclusively using normal log data for masked token prediction. The threshold for discerning deviations in new log data from established normal patterns was determined based on the accuracy of top-k token prediction.
In the vehicular domain, a two-stage approach for detecting intrusions in the Controller Area Network (CAN) was proposed [15]. The approach involves pre-training BERT on normal CAN data to understand its semantic representation, followed by fine-tuning the model to classify messages as either normal or malicious. Despite its promising results, the model evaluates each packet individually, overlooking the sequential nature of some attacks. In contrast, in [26], a sliding window of CAN-IDs is fed to BERT to detect message injection in the CAN-bus. The model is trained using masked language self-supervised training to predict a set of candidates CAN IDs for a masked ID. If the actual ID does not correspond to one of the predicted IDs, it is considered abnormal. The results indicate that BERT-based approaches outperform other ML methods such as PCA and Autoencoders in terms of anomaly detection performance. In another work [27] the authors explore the use of well-known LLMs for CAN intrusion detection. They examine different architectures where an LLM is combined with a classification head, utilizing models like SecureBERT, BERT, and LLAMA-2. These models are fine-tuned on a database related to CAN. The results confirm the effectiveness of the models and highlight the superiority of larger models such as LLAMA-2 in capturing intricate message distributions.
In addition to research concentrating on in-vehicle communication security, a transformer was utilized for detecting misbehavior in V2X communication, as discussed in [14]. The authors fine-tuned an encoder-based transformer using normal BSM sequences, each with a length of 200, to replicate them. Subsequently, misbehavior detection was accomplished by comparing the differences between a misbehaving vehicle’s sequence and its transformed counterpart against a predefined threshold. While the results are promising, processing a 200 BSM input sequence may pose computational limitations, as can behavior evaluation with an insufficient number of BSMs. Another work using transformers is presented in [28], where a privacy-preserving communication architecture is proposed, utilizing a transformer-based Intrusion Detection System (IDS) trained on edge devices with data from connected vehicles. The cloud aggregates the model weights to distribute the final global model. The IDS utilizes a tailored encoder-based transformer, which integrates feature attention and routing, to choose the best feature sets from network packets for classification. Despite the promising results, a limitation of this work arises from the choice of a packet-wise approach, disregarding the importance of the packet context for the detection. In addition, conducting tests on non-vehicular data raises questions regarding the model’s effectiveness in detecting specific attacks on vehicular networks.
III Background
III-A Vehicular communication models
V2X communications encompass a variety of technologies facilitating real-time information exchange between vehicles and their surroundings, thereby enhancing road safety, and traffic efficiency. Figure 1 provides an overview of the different types.
V2V communication involves vehicles equipped with network modules, called onboard units, enabling them to exchange data concerning their position, speed, heading, and acceleration. This type of communication aims to prevent accidents by facilitating situational awareness among vehicles.
In contrast, V2I communication allows vehicles to interact with roadside units (RSUs) such as streetlamps, RFID readers, lane markers, and signage. This bidirectional communication enables the real-time exchange of information about road conditions, traffic, accidents, and more, thereby enhancing overall traffic management and safety[29].
V2P, or vehicle-to-pedestrian communication, involves interactions with vulnerable entities in the environment, such as pedestrians and cyclists. These individuals typically carry smartphones or wearable devices capable of transmitting their location to alert nearby vehicles of their presence, contributing to enhanced safety for all road users.
Additionally, V2X includes V2N communication, which enables vehicles to connect to cloud-based services and platforms. This connectivity facilitates access to a wide range of services such as over-the-air updates, remote diagnostics, real-time traffic and navigation information, and data analytics for traffic management purposes.
In order to protect ensure vehicle privacy, vehicles use pseudonyms for communications. Vehicle pseudonyms are temporary, randomly generated identifiers used in vehicular communication systems to provide privacy and anonymity for vehicles while still allowing them to communicate with each other and infrastructure. These pseudonyms are periodically changed to prevent tracking of vehicles over time.
Figure 1: V2X communication and Misbehavior detection with edge and cloud LLMs
III-B Transformers and LLMs
The transformer is a groundbreaking architecture that revolutionized the task of sequenced data processing, introducing innovative mechanisms for capturing complex dependencies and patterns within sequences. This was achieved by leveraging the self-attention mechanism to weight the importance of each token in sequence with regard to the other tokens. The weights are obtained by applying three linear transformations, namely (keys (k), queries (Q) and values (V)) on the input sequence of size $d_{k}$ following the equation 1.
Several variations of transformers have been proposed and customized to address a broad spectrum of tasks, encompassing language understanding, generation, translation, and more. Notable examples include the original Transformer architecture [30] which is presented in Figure 2, comprised of both encoder and decoder blocks, as well as BERT [12] and RoBERTa [31], both of which are encoder-only transformers. Additionally, LLAMA [11] and GPT exemplify the decoder-only type of transformers, primarily designed for generation tasks.
Mistral-7B, is also a decoder-only transformers that employs specific design decisions, allowing it to outperform models such as the 13 billion-parameter variation of Llama-2 and the 34 billion-parameter Llama-1 with only 7 billion parameters. Those technical choices involve using Sliding Window Attention (SWA) that operates by defining a fixed-size window surrounding each token, thus limiting the number of tokens it has to tend for [32]. Compared to the vanilla self-attention used in [30], this technique enables longer sequence processing with lower computation cost. Moreover, Mistral-7B employs the Grouped-query attention (GQA) technique where a group of queries shares a single key and value heads[33], thus speeding up the inference time of the model.
[…]
Figure 2: Transformer architecture
III-C Low rank adaptation
In traditional fine-tuning, the model’s weights are directly modified to adapt the model to the new considered task. While fine-tuning is a crucial approach ensuring transfer learning, the process become computationally daunting as the number of parameter increases. Notably, fine-tuning large language models poses significant challenges in terms of memory requirements and computational resources.
Low Rank Adaptation (LoRA), is a technique that was proposed in [34] to alleviate the memory footprint and speed up the process of large models’ fine-tuning. This is achieved through reducing the number of trainable parameters by decomposing the changes on the pre-trained weight matrices into low rank matrices. The process suppose that after fine-tuning the new weights matrices will have the form $W=W+\Delta W$ where W is frozen during training and $\Delta W$ is decomposed into two lower rank matrices A and B. Being of lower rank, the number of trainable parameters in these matrices is smaller than the original $\Delta W$.
Although LoRA enhances significantly the computational requirements, further optimization could be achieved by using QloRA [35], which applies quantization techniques to further reduce the memory footprint.
IV System model
We present an approach centered around the utilization of Mistral-7B for detecting vehicle misbehavior. Our method involves reframing the misbehavior detection problem as a sequence classification task. In this framework, we deploy a model based on Mistral-7B, which is quantized and implemented on a roadside unit (RSU). Through quantization, the memory requirements for running Mistral-7B are significantly reduced, enabling its deployment on edge equipment at the RSUs. These RSUs possess the necessary capabilities to collect BSMs exchanged within their sensing ranges and execute the model to identify potential misbehavior.
The collected messages undergo the prepossessing steps presented in Figure 3-a, where the irrelevant features are deleted, the messages are grouped by sender pseudonym and then transformed into BSM-Windows of size $n$ for each user. Following the prepossessing steps, the formed time windows are transformed into textual prompts necessary for the language model inference. The prompt design steps transform the multivariate time window to a text, as described in Figure 3-b. The obtained prompt is then fed to a Tokenizer and then Mistral-7B-based classifier for the final labeling. The last step following the behavior assessment is the reporting phase, where the RSU communicates the information’s on the misbehaving vehicle to a trust authority (TA) located in the cloud. The TA is then responsible for verifying the report, and feeding it, if necessary, to the cloud LLM, which can complement the report by conducting deeper analysis and validation, and incorporating additional information from a broader dataset stored in the cloud. Subsequently, the TA takes the appropriate response that could include revoking the vehicle’s access to the network.
It is important to highlight that our methodology ensures the examination of both contextual and content aspects within the messages at the edge LLM. Contextual dimensions are considered during the grouping of messages into time windows, facilitating the analysis of temporal patterns and relationships. Simultaneously, the content aspect is considered by a detailed inspection of the message contents.
Figure 3: Proposed architecture
V Experiments and Results
In this section, we delve into the details of the conducted experiments. We start by presenting the preprocessing steps we applied on the database used, namely the VeReMi dataset. Thereafter, we define the performances metrics used to evaluate the performance of our model. After this, we present the two major sets of experiments, binary and multiclass classification. For each set, we investigate two possible methods of leveraging Mistral-7B: one involving fine-tuning and the other without. Our aim is to assess whether we can leverage Mistral-7B with minimal modifications, maximizing its utility without extensive adjustments. From there, we compare its performance against two other LLMs, LLAMA2-7B and roBERTa. Lastly, we analyze the effect of varying window sizes on the inference time to identify the necessary requirements and potential solutions for deploying the models.
V-A Extended VeReMi dataset
VeReMi [36] is a Vehicular Reference of Misbehavior dataset that was created to evaluate the performance of misbehavior detection systems in the VANETs. It was generated using the Framework For Misbehavior Detection (F2MD), which is an extension of the vehicular network simulator VEINS, and a subsection of the Luxembourg SUMO Traffic (LuST) scenario covering 1.61 km² with a peak density of 67.4 vehicles per km². To experiment and assess our approach, we used the extended version of VeReMi[37] that offers more diverse attack types including DoS, Random Positioning Attacks, Constant Positioning Attacks, Disruptive Attacks, Constant Speed Attacks, Delayed Messages Attack, Data Replay Attacks, Traffic Sybil Attacks, and Eventual Stop Attacks that were simulated under varying traffic conditions specifically from 07 to 09 am and from 2 to 4 pm. The dataset is structured into trace files containing the exchange of BSMs by individual vehicles. During each attack simulation, vehicle behaviors are stored in separate files encompassing various fields including: type, receive time, sending time, sender ID, sender Pseudo, message ID, position, position noise, speed, speed noise, acceleration, acceleration noise, heading, and heading noise.
Preprocessing of the raw files involved extracting all messages, concatenating them, and selecting only type-3 messages, which describe the received BSMs. Duplicate entries resulting from the interception of the same message by multiple vehicles were eliminated. Then fields describing noise attributes such as position noise, heading noise, acceleration noise and speed noise were dropped. In order to minimize memory usage during fine-tuning and inference, adjustments were made to the structure of certain fields. The heading, position, acceleration, and speed fields, originally consisting of lists of three values each representing the x, y, and z coordinates, underwent transformation. For speed and acceleration, the norm was calculated, condensing the data into single values. Additionally, the position and heading fields were split into two distinct fields: position-x and position-y, and heading-x and heading-y, respectively. Given that the z coordinate remained consistently null across all entries, it was removed. Lastly, to further reduce memory requirements, all numerical values were rounded to three decimal points. Following this, each message was labeled based on its sender ID, therefore, every message issued by a misbehaving car was labeled as attack. This classification is derived from the title of the corresponding trace file, referencing the attack-code correspondence outlined in Table I. Once the dataset was labeled, messages for each sender ID were grouped, ordered by the receive time and segmented into time windows.
The final preprocessing step, is the textualization of the time windows, which could be achieved either by considering each row as a separate sentence having the feature values preceded by the corresponding field name, or column wise like described in Fig. 3-b. As the former results in large prompt size due to repeated field names, we resorted to the column-wise approach.
For our experiments, We selected a subset of attacks from the VeReMi dataset, and we processed them to extract time windows. In order to be able to select what is the best number of packets inside a time window, we created different datasets with 4 sizes, 10, 20, 50 and 100. Table I and Figure 4 give an overview of the selected attacks and class distribution after processing.
Identifier Attack Type Attack description
A0 Genuine Normal behaving vehicle with no malicious intention
A1 ConstPos Vehicle transmitting constant position attributes
A5 ConstSpeed Vehicle transmitting constant speed field.
A9 EventualStop
Vehicle simulating a sadden stop by freezing its position and setting
its speed to zero.
A11 DataReply Vehicle retransmitting messages received from a target vehicle.
A12 DelayedMessages Vehicle transmitting delayed correct information.
A13 DoS Vehicle transmitting at rate exceeding the predefined limit.
A14 DoSRandom Malicious V2X node, conducting DoS with transmitted fields set randomly.
A18 DoSRandomSybil Malicious node would use Sybil attack to conduct DoS random.
TABLE I: Mapping attack type to codes[37]
Figure 4: The class distribution after dataset preprocessing
V-B Performance Metrics
To assess the performance of our misbehavior detector, we used the following metric suit.
Recall
measures the proportion of actual positives that are correctly identified by the system. A high recall reflects the detector’s ability to identify the actual attack and minimize the chances of labeling an attack as normal.
Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ (2)
Precision
Calculates the correctly labeled attacks to the actual attacks, forming an idea on the model’s capacity to identify attacks while minimizing false alarms.
Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$ (3)
F1 score
It is the harmonic mean average of the recall and precision, allowing to assess the model’s capacity to balance the two, thus providing a balanced evaluation of the system’s performance.
$F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (4)
Accuracy
Computes the percentage of instances correctly classified by the model. It reflects the model’s ability to correctly identify the normal from the anomalous behavior.
Accuracy=True Positives+True NegativesTotal $\text{Accuracy}=\frac{\text{True Positives}+\text{True Negatives}}{\text{TotalInstances}}$ (5)
V-C Binary classification
In the first set of experiments, we utilize Mistal-7B to distinguish between normal behaving cars and misbehaving ones. In the following subsections, we outline the different settings we experimented with Mistral-7B, and we compare its performance against two other renown LLMs.
V-C1 Fine-tuned Mistral
We employed qLoRA for fine-tuning Mistral-7B, leveraging the hyperparameters outlined in Table II. The classifier is formed by Mistral-7B transformer, on top of which a linear classification head based on the last token is added.
We chose to perform the fine-tuning on the self-attention modules and the weights of the classification head to maximize performance while minimizing parameter count, as recommended in [34]. Furthermore, the $r$ and the LoRA alpha in the Table II refer to the dimension of the low-rank matrices and the scaling factor which calibrates the weights update magnitude during the training, respectively. Figure 5 presents the effect of the hyperparameter r on the accuracy when window size equals 10. The figure shows that the best accuracy is obtained for an r value of to 2, whereas it decreases significantly for higher values. This justifies the choice indicated in Table II.
Figure 5: The effect of low-rank matrix rank on the evaluation accuracy.
The findings presented in Table III summarize the outcomes of fine-tuning Mistral-7B. We can notice that, the more the window size increases, the better the model is able to classify vehicle behavior. This is because larger windows offer more information on the messages communicated by the vehicle, making the classification task easier. The most favorable results were achieved using a window size of 100, yielding an accuracy and F1 scores of 98%. Notably, this was achieved despite the utilization of only 1000 instances per class and tuning only 0.012% of the parameters.
Hyperparameter Values
Quantization type 4-bit
$r$ 2
LoRA alpha 16
LoRA dropout 0.05
Bias none
Target modules [q_proj, v_proj, score]
Task type sequence classification
TABLE II: Hyperparameter settings for fine-tuning with QLoRA
Window size Accuracy F1 Recall Precision
10 0.945 0.944933 0.945445 0.948269
20 0.970 0.969988 0.970297 0.971429
50 0.980 0.979998 0.980198 0.980583
100 0.980 0.98020 0.980120 0.98028
TABLE III: Binary Mistral-7B results
V-C2 Mistral-7B as feature extractor
For this set of experiments, we train a classification neuron network on top of frozen Mistral-7B layers, that we call MistralFeatureNet. The purpose of MistralFeatureNet is to study the capacity of the transformer to extract relevant features from the textualized windows without any fine-tuning. The classification process starts by collecting Mistral-7B’s top layer output of shape $(n_{tokens},4096)$, then max-aggregating it into a 4096-sized vector. Thereafter, the resulting vector is passed through several dense layers for the final classification. The number of hidden layers and their specifications are indicated in Table IV.
We performed the training on a balanced basis between normal and abnormal cases. Since size 100 contains the fewest data points, we were careful to select a similar class distribution for the other sizes, i.e. 21923 per class. Table V summarizes the results. The table shows the model’s positive outcomes, especially with 10 and 20 window sizes, with an accuracy and F1 above 90%. In addition, we observe that for bigger window sizes, the performance deteriorates. A possible cause of this is the max aggregation step. Since we are taking the maximum over the tokens’ representation on each model dimension, the larger the sequence, the higher the chance of loosing information.
Hyperparameter Value
Input dimension 4096
Hidden Layers dimension [256, 128, 64, 32]
dropout rate 0.2
Optimizer Adam
Learning rate 0.0001
Number of epochs 80
Batch size 32
TABLE IV: Binary classification head hyperparameters
Window size Accuracy F1 Recall Precision
10 0.94100 0.940000 0.902000 0.980000
20 0.94200 0.943000 0.930000 0.980000
50 0.93000 0.93000 0.80000 0.990000
100 0.820000 0.900000 0.860000 0.85000
TABLE V: Binary MistralFeatureNet results
V-C3 Other LLMs for binary classification
To further validate the model choice, we fine-tuned roBERTa and LLAMA2-7B under similar conditions. We employed similar qLoRA parameter described in TableII. With LLAMA2-7B we obtained good results with only 1000 instance per class, whereas roBERTa required a minimum of 3000 per class to converge. Figure 6 summarizes the different models performance across the time windows. The immediate observation is the superiority of fine-tuned Mistral-7B over all the other models on most metrics, notably the accuracy and the F1 score. Conversely, RoBERTa, yields the worst results for all metrics. This is due to it being the only model with a predefined maximum sequence length of 512, i.e., roBERTa cannot process token sequences of size surpassing 512. The most common solution to circumvent this limitation is to truncate the sequence to fit the model’s requirements. However, most of our prompts surpass this value. The effect of the truncation is moderate for window size 10, however it becomes more prominent as the size of the prompts grows. Another contributing factor to this behavior is the textualization step described in Section V-A, in which each column is considered as a sentence, so the truncation step results in eliminating whole features from the sequence.
Finally, Figure 6 suggests that 100 is the most suitable window size for the fine-tuned Mistral-7B. In contrast, 50 is the most suitable for LLAMA2-7B and MistralFeatureNet as they exhibit a deterioration when the window size is increased further.
Figure 6: Exploring Classification Metrics Variability among Binary Models
V-D Multiclass classification
For multi-class classification, the goal is to use Mistral-7B to detect and correctly classify attacks among the 9 classes described in Table I. We employ the identical reasoning as in the binary categorization. We fine-tune Mistral-7B, and then we compare its results against, MistralFeatureNet for the multi-class classification, LLAMA2-7B and roBERTa.
V-D1 Fine-tuned Mistral-7B
Similarly to the binary classification with Mistral-7B, we leverage qLoRA fine-tunning with similar hyperparameters, as they allow minimizing the computational cost. In this case, we select 1000 samples of each class to ensure the balance between the classes.
Figure 8 shows the performance results for different window sizes. As with binary classification, the hypothesis that the model performs better with a larger window is validated. The refined model obtained an F1 score of 0.96 in a 9-class classification task with just 1,000 instances per class. The confusion matrix in Figure 7 also proves the potential of the model, which correctly discerns most classes.
Figure 7: Mistral-7B Confusion Metrics
V-D2 Mistral for feature extraction
Following the procedure of the binary classification, we freeze the layers of Mistral-7B, and we train a neuron network classifier leveraging the hyperparameters presented in Table VI. Since the task involves a considerable number of classes, the convergence of the model required a bigger training set, thus we employed all the dataset. However, our data is unbalanced with a lot more benign instances, therefore, we undersampled the genuine class to have a comparable representation to other classes.
The results, presented in Table VII, reflect the challenges encountered by the model in the 9-class classification task. The best outcome was scored with a window size of 20. However, the performance significantly drops for bigger window sizes. Therefore, for multiclass classification, fine-tuning is the best option to obtain good results. This is why we retained the fine-tuned.
Hyperparamter Value
Input dimension 4096
Hidden Layers dimension [256, 128, 64, 32]
dropout rate 0.2
Optimizer Adam
Learning rate 0.0001
Number of epochs 150
Batch size 64
TABLE VI: Classification network hyperparameters
Window Accuracy F1 Score Recall Precision
10 0.8692 0.8486 0.8692 0.8461
20 0.8758 0.8537 0.8758 0.8490
50 0.8385 0.8122 0.8385 0.8164
100 0.6128 0.5654 0.6128 0.5587
TABLE VII: Multi-class MistralFeatureNet results
V-D3 Other LLMs for multiclass classification
The performance of LLAMA2-7B and roBERTa on the nine-class classification task is illustrated in Figure 8. During training, LLAMA2-7B was configured similarly to Mistral-7B, with each class containing 1000 instances. In contrast, roBERTa required a minimum of 4000 instances per class. The results highlight the superiority of large Transformers’ over smaller ones. It is observed that LLAMA-7B achieves a similar accuracy to Mistral-7B, both around 0.95. Conversely, roBERTa consistently exhibits the least beneficial outcomes. Finally, we find that larger models, particularly fine-tuned Mistral-7B, exhibit superior processing capabilities for complex data messages and consistently outperform smaller LLMs such as roBERTa in the multiple-class classification task. Moreover, they achieve this capacity with a reduced number of samples, a crucial advantage when fewer vehicle communication data samples are available at the edge, which can occur due to fluctuations in traffic.
Figure 8: Exploring Classification Metrics Variability among multiclass classifiers
V-E Inference time and deployment considerations
Due to the size of the Mistral-7B model, deploying it requires a balance between multiple hyperparameters to avoid adding overhead to a process that requires speed and accuracy. One method to mitigate memory demands is through quantization. In our experiments we used 4-bit quantization, which reduces significantly the memory required to load the model to 4GB.
In addition to memory requirements, it is also very important to study the inference latency. It is evident that the size of the prompt is one of the main factors affecting inference time. While we demonstrate that the best window size for binary and multiclass classification is 100 using fine-tuned Mistral-7B, opting for these values might introduce inference latency which is limitating to the global performance of the MDS. Figure 9, illustrates the evolution of the inference time with regard to the window size for Mistral-7B and LLAMA2-7B. We observe similar trends for both models, where the inference time increases notably as the window-size increases. This patern may pose a constraint on the use of LLMs in this context. Importantly, these measurements were conducted on A100 40GB GPUs and depending on the hardware, actual values may vary from the reported figures.
However, these models are, designed to seamlessly process prompts with different sizes without having to modify the architecture or apply additional preprocessing, which guarantees flexibility in terms of the window-size. Therefore, based on the computational resources available, the size can be adapted to enhance performance without notable accuracy reduction. the figures have shown that Mistral-7B. maintains good performance even with smaller window sizes.
Figure 9: Inference Time per Window Size
VI Conclusion
In this paper, we explored the use of Large Language Models for vehicle misbehavior detection in vehicular networks, particularly by addressing the challenges of using LLMs at the edge, as part of a cloud-edge LLM based MDS system. We leveraged the new state-of-the-art LLM, Mistral-7B, to detect abnormal behaving vehicles. Our detection process involves transforming a sequence of features extracted from BSMs into a text prompt, which is then processed by a sequence classification mechanism using Mistral-7B. Through multiple experiments, we have demonstrated the efficacy of fine-tuning Mistral-7B using QLoRA, even with a limited dataset, across both binary and multiclass misbehavior classification tasks. Furthermore, our comparative analysis revealed the superior performance of Mistral-7B over concurrent LLMs, including LLAMA2-7B and roBERTa. Subsequently, we investigated the optimal number of examined packets or ”window size”, in a prompt to achieve the best results. Our finding show that larger window sizes tends to enhance misbehavior classification.
In addition to evaluating model performance, we explored the practical considerations of deployment. We have addressed deployment challenges and proposed strategies such as quantization and the selection of an adequate window size to accommodate hardware constraints. Notably, our model’s intrinsic flexibility with regard to varying window sizes, derived from its use of a language model, allows its adaptability to constraints of hardware configurations. In conclusion, our study offers promising results as well as valuable practical insights for deploying misbehavior detection systems utilizing LLMs within vehicular networks contexts.

### v2

@@
-MistralBSM: Leveraging Mistral-7B for Vehicular Networks Misbehavior Detection
+MistralBSM: Leveraging The Mistral-7B LLM for Vehicular Networks Misbehavior Detection
 Abstract
-Vehicular networks are exposed to various threats resulting from malicious attacks.
-These threats compromise the security and reliability of communications among road users, thereby jeopardizing road and traffic safety.
-One of the main vectors of these attacks within vehicular networks is misbehaving vehicles.
-To address this challenge, we propose deploying a pretrained Large Language Model (LLM)-empowered Misbehavior Detection System (MDS) within an edge-cloud detection framework.
-Specifically, we fine-tune Mistral-7B, a state-of-the-art LLM, as the edge component to enable real-time detection, whereas a larger LLM deployed in the cloud can conduct a more comprehensive analysis.
-Our experiments conducted on the extended VeReMi dataset demonstrate Mistral-7B’s superior performance, achieving 98% accuracy compared to other LLMs such as LLAMA2-7B and RoBERTa.
-Additionally, we investigate the impact of window size on computational costs to optimize deployment efficiency.
-Leveraging LLMs in MDS shows interesting results in improving the detection of vehicle misbehavior, consequently strengthening vehicular network security to ensure the safety of road users.
-Index Terms:
-Vehicular network, Large Language Models (LLMs), misbehavior detection, intrusion detection
+Malicious attacks on vehicular networks pose a serious threat to road safety as well as communication reliability.
+A major source of these threats stems from misbehaving vehicles within the network.
+To address this challenge, we propose a Large Language Model (LLM)-empowered Misbehavior Detection System (MDS) within an edge-cloud detection framework.
+Specifically, we fine-tune Mistral-7B, a compact and high-performing LLM, to detect misbehavior based on Basic Safety Messages (BSM) sequences as the edge component for real-time detection, while a larger LLM deployed in the cloud validates and reinforces the edge model’s detection through a more comprehensive analysis.
+By updating only 0.012% of the model parameters, our model, which we named MistralBSM, achieves 98% accuracy in binary classification and 96% in multiclass classification on a selected set of attacks from VeReMi dataset, outperforming LLAMA2-7B and RoBERTa.
+Our results validate the potential of LLMs in MDS, showing a significant promise in strengthening vehicular network security to better ensure the safety of road users.
 I Introduction
-The evolution of vehicular networks [1], commonly known as vehicle-to-everything (V2X), has been marked by continuous developments in communication and transportation technologies.
-Originally conceived as a means of improving vehicle safety through the exchange of information between vehicles and the infrastructure, research expanded into areas such as cooperative driving, advanced scenarios to support automated driving, and the integration of vehicle networks with emerging technologies such as 5G and beyond [2] and the Internet of Things (IoT) [3].
-Today, as the potential of connected autonomous vehicles promises to redefine the future of private and public transportation, the evolution of vehicular networks research continues apace, particularly with regard to cybersecurity issues[4]
+The evolution of vehicular networks [1], commonly known as vehicle-to-everything (V2X), has been marked by continuous developments in communication, transportation, and consumer technologies.
+Originally conceived as a means of improving vehicle safety through the exchange of information between vehicles and the infrastructure, research expanded to include more advanced scenarios such as cooperative, automated driving and more consumer-centric applications such as in-vehicle infotainment systems, personalized mobility experiences, and seamless connectivity with consumer personal devices [2].
+This evolution has been accelerated by the integration of emerging technologies like 5G and beyond [3] and the Internet of Things (IoT) [4], which enable more sophisticated, consumer-oriented services.
+Today, as the potential of connected autonomous vehicles promises to reshape the landscape of personal and public transportation, the evolution of vehicular networks research continues apace, focusing on enhancing consumer experiences.
+This includes addressing network security challenges [5], enabling a reliable and secure experience within the intelligent transportation scene.
 Indeed, vehicular networks are exposed to various threats resulting from malicious attacks [6].
 These threats compromise the security and reliability of communications among road users, thereby jeopardizing road and traffic safety.
-One of the main vectors of these attacks within vehicular networks is misbehaving vehicles [6].
-Misbehaving vehicles can launch denial-of-service (DoS) attacks which flood the network with malicious traffic, disrupting communication between vehicles [7].
+One of the main vectors of these attacks within vehicular networks is vehicle misbehavior [7].
+Misbehaving vehicles can launch denial-of-service (DoS) attacks, which flood the network with malicious traffic, disrupting communication between vehicles [8].
 As a result, this disruption can lead to communication breakdowns, making vehicles unaware of each other and disrupting the operation of both traffic systems and vehicular safety applications.
@@
 Vehicles interact unpredictably, leading to various behaviors that may differ from expected patterns.
-Additionally, the large volume of data generated by vehicles adds complexity.
+Additionally, the large volume of data generated by vehicles increases the complexity.
 Analyzing this data in real-time requires advanced algorithms and computing resources.
-Machine learning (ML) has recently emerged as an interesting technology in the development of misbehavior detection systems (MDS).
-ML enables MDS to adapt and learn from historical data, continuously improving its ability to identify abnormal or malicious behavior.
-[9].
+Machine learning (ML) has recently emerged as an interesting solution in the development of Misbehavior Detection Systems (MDS).
+ML enables MDS to adapt and learn from historical data, continuously improving its ability to identify abnormal or malicious behavior [10].
 With the recent advancements in large language models (LLMs) there is an exciting opportunity to further bolster the capabilities of MDS.
-Indeed, models, such as the GPTs [10], LLAMA-2[11] and BERT[12] have demonstrated remarkable capabilities in various natural language processing tasks, but also showed interesting capacities in network security tasks [13], including for vehicular networks security [14, 15].
+Indeed, models such as the GPTs [11], LLAMA-2[12] and BERT [13] have demonstrated remarkable capabilities in various natural language processing tasks but also showed interesting capacities in network security tasks [14, 15, 16], including for vehicular networks’ security [17, 18].
 The potential of LLMs in vehicle misbehavior detection stems from the fact that LLMs can facilitate the extraction of insights from textual data of vehicle communications.
-Through semantic analysis and contextual understanding, these models can identify linguistic features that can signify potential abnormal behavior patterns.
+Through semantic analysis and contextual understanding, these models can identify relevent features that can signify potential abnormal behavior patterns.
 However, using LLMs for vehicle misbehavior detection poses significant challenges, notably in privacy and resource constraints.
@@
 Nevertheless, transferring all vehicle communication data to cloud infrastructure poses potential risks for real-time misbehavior detection and raises concerns about data privacy.
-Recently, a new large-scale model named Mistal-7B [16] has been introduced, rivaling existing models with fewer parameters thanks to various architectural choices.
-This opens the possibility to use an LLM-empowered MDS architecture with edge-cloud collaboration, where a light LLM Model such as Mistral-7B resides at the edge (e.g., at a roadside unit), and a bigger LLM resides in the cloud.
-Edge LLMs can perform initial vehicle misbehavior detection in real-time, providing immediate feedback and alerts to vehicle systems.
-Cloud LLMs can complement this by conducting deeper analysis, incorporating additional contextual information from a broader, and possibly multi-modal dataset stored in the cloud (e.g., multimodal traffic information from the intelligent transportation system).
+Recently, several LLMs were introduced, among which Mistral-7B [19] has gained attention for its exceptional balance between model size and performance, making it particularly suitable for ressource constrainted application like vehicular networks.
+More importantly, it is open-weight model, which opens the possibility to adapt and leverage it as an LLM-empowered MDS architecture.
+In this work we propose an edge-cloud setting, where a quantized Mistral-7B is deployed at the edge (e.g., at a roadside unit (RSU), and a deeper LLM resides in the cloud.
+The Edge LLM can perform initial vehicle misbehavior detection in real-time on short Basic Safety Message (BSM) sequences, providing immediate feedback and alerts to vehicle systems.
+Cloud LLMs can complement this by conducting more in-depth analysis on longer BSM sequences and by incorporating additional contextual information from a broader and possibly multi-modal dataset stored in the cloud (e.g., multimodal traffic information from the intelligent transportation system).
 Using the edge LLM has the advantage of processing vehicular communication data locally, minimizing the need to transmit raw data to the cloud and preserving user privacy.
-Cloud LLMs can be utilized for more general analysis that uses edge LLMs insights, and other data external to the communication system, ensuring compliance with privacy regulations and minimizing the risk of data privacy breaches.
-In this article, we explore the use of Mistral-7B as an edge LLM to be located at a roadside unit, and we explore the capacity of the LLM to classify each vehicle as normal or misbehaving.
-The main contributions of this article are as follows:
-  • We use Mistral-7B for misbehavior detection based on the analysis of observed basic safety massages (BSMs).
-  • We transpose a vehicular misbehavior detection task into a sequence classification by converting windows of BSMs originating from a specific vehicle into text prompts to be classified by a language model.
-  • We fine-tune Mistral-7B as a sequence classifier to assess a vehicle behavior nature through a sequence of its broadcasted BSMs, resulting in MistralBSM.
-  • We assess Mistral-7B’s performance compared to other prominent open-source LLMs, namely LLAMA2-7B and BERT based architecture, in the context of MDS, and we demonstrate its superior results.
+Cloud LLMs can be utilized for more general analysis that uses edge LLMs insights and other data external to the communication system, ensuring compliance with privacy regulations and minimizing the risk of data privacy breaches.
+In this article, we focus mainly on the implementation of Mistral-7B as the edge LLM to classify vehicle behaviors. Thus, the main contributions of this article are as follows:
+  • We propose an LLM-based Misbehavior detection framework leveraging Mistral-7B for BSM sequence classification.
+  • We design our data preprocessing carefully to avoid label leaking or training data leaking for a more reliable performance assessment and efficient model inference to enable easy edge deployment.
+  • We compare Mistral-7B’s performance to other prominent open-source LLMs, namely LLAMA2-7B and BERT-based architecture, in the context of MDS, and we demonstrate its superior results.
+  • We use quantization and study BSM sequence lengths to optimize model deployment within the limited computational resources of vehicular networks.
 The rest of this paper is organized as follows.
 Section II introduces the related works of misbehavior detection methods in vehicular networks.
-Section III introduces a background about vehicular networks architecture, LLMs and low rank adaptation for LLM fine-tuning.
-In Section IV, we define our system model, the proposed solution and introduce Mistral for misbehavior detection.
+Section III introduces a background about vehicular networks’ architecture, LLMs and low-rank adaptation for LLM fine-tuning.
+In Section IV, we define our system model and introduce Mistral for misbehavior detection.
 In Section V, we do a performance evaluation of our model and compare our results.
 II Related works
-II-A Vehicular misbehavior detection
-Detecting misbehavior is widely recognized as a critical task in V2X security.
-ML-based approaches are commonly favored due to their distinctive ability to model and classify misbehavior, surpassing other rule-based strategies.
-In [17], various models including Random Forest, Decision Tree, Ada Boost Classifier, Ridge Classifier, Naive Bayes, and XGBoost have been evaluated on the CAN-OITDS dataset for multi-class classification involving four distinct attack types.
-The results showed an advantage of XGBoost.
-The authors of [18], also employed ML technique for misbehavior detection.
-The proposed framework enhances the detection performance by including movement and position plausibilities with the BSM features in the input of the classifier.
-The considered algorithm includes SVM, KNN, Naive Bayes, Random Forest, ensemble boosting and voting ensemble.
-The primary drawback of such models is their limited ability to adequately handle the temporal aspect of specific attacks aimed at V2X communication.
-To address this gap, some researchers considered using LSTM based architectures.
-For example, in [19], [20] and [21], LSTM and CNN were combined to detect misbehavior, yielding promising results.
-On the other hand, in [22] and [23], considered the misbehavior detection as time-series anomaly detection problem that could be modeled by a Markov Decision process where each decision is based on the previous ones.
-They proposed a DQN-based classifier where the agent q-network is formed by LSTM layers and a fully connected layer.
-Particularly, in [22], the proposed framework uses KNN as an unsupervised algorithm for labelling BSMs.
-Thereafter, the labeled data is used by the DQN agent to learn the optimal policy.
-The effectiveness of these techniques underscores the importance of considering the temporal dimension in detecting misbehavior within vehicular networks, where LSTM has demonstrated efficacy.
-However, interest has shifted with the introduction of transformers and LLMs, as they enable efficient processing of sequenced data.
-II-B LLMs for network security
-LLMs have showcased remarkable capabilities that extend beyond natural language processing (NLP), generating considerable interest across diverse domains.
-Notably, they have gained interest in network security, with encoder-based transformers being particularly renowned.
-For example, BERT served as the foundation for log anomaly detection systems as described in ([24], [13], [25]).
-In [24], BERT was trained on normal log key sequences extracted from raw network logs to predict the next masked key.
-Subsequently, during the detection phase, the model predicted normal keys with high precision, while abnormal keys exhibited low detection accuracy.
-Conversely, in [13], the authors fine-tuned BERT using a fully connected classifier, where the input consists of a time window comprising a sequence of parsed log data.
-In [25], log anomaly detection is achieved by fine-tuning LongFormer, a transformer based on BERT, exclusively using normal log data for masked token prediction.
-The threshold for discerning deviations in new log data from established normal patterns was determined based on the accuracy of top-k token prediction.
-In the vehicular domain, a two-stage approach for detecting intrusions in the Controller Area Network (CAN) was proposed [15].
-The approach involves pre-training BERT on normal CAN data to understand its semantic representation, followed by fine-tuning the model to classify messages as either normal or malicious.
-Despite its promising results, the model evaluates each packet individually, overlooking the sequential nature of some attacks.
-In contrast, in [26], a sliding window of CAN-IDs is fed to BERT to detect message injection in the CAN-bus.
-The model is trained using masked language self-supervised training to predict a set of candidates CAN IDs for a masked ID.
-If the actual ID does not correspond to one of the predicted IDs, it is considered abnormal.
-The results indicate that BERT-based approaches outperform other ML methods such as PCA and Autoencoders in terms of anomaly detection performance.
-In another work [27] the authors explore the use of well-known LLMs for CAN intrusion detection.
-They examine different architectures where an LLM is combined with a classification head, utilizing models like SecureBERT, BERT, and LLAMA-2.
-These models are fine-tuned on a database related to CAN.
-The results confirm the effectiveness of the models and highlight the superiority of larger models such as LLAMA-2 in capturing intricate message distributions.
-In addition to research concentrating on in-vehicle communication security, a transformer was utilized for detecting misbehavior in V2X communication, as discussed in [14].
-The authors fine-tuned an encoder-based transformer using normal BSM sequences, each with a length of 200, to replicate them.
-Subsequently, misbehavior detection was accomplished by comparing the differences between a misbehaving vehicle’s sequence and its transformed counterpart against a predefined threshold.
-While the results are promising, processing a 200 BSM input sequence may pose computational limitations, as can behavior evaluation with an insufficient number of BSMs.
-Another work using transformers is presented in [28], where a privacy-preserving communication architecture is proposed, utilizing a transformer-based Intrusion Detection System (IDS) trained on edge devices with data from connected vehicles.
-The cloud aggregates the model weights to distribute the final global model.
-The IDS utilizes a tailored encoder-based transformer, which integrates feature attention and routing, to choose the best feature sets from network packets for classification.
-Despite the promising results, a limitation of this work arises from the choice of a packet-wise approach, disregarding the importance of the packet context for the detection.
-In addition, conducting tests on non-vehicular data raises questions regarding the model’s effectiveness in detecting specific attacks on vehicular networks.
+Misbehavior detection is a pivotal component in vehicular networks where ML can offer an advantage over rule-based approaches given the significant amount of communication data generated by the network and required to achieve an accurate detection.
+In the literature, several ML-approaches have been proposed, from simple algorithms to deep models, to address in-vehicle attacks as well as V2X communications misbehavior.
+For instance, the authors [20] studied several famous ML algorithms, including tree-based models such as Decision Tree, XGBoost, and Ada Boost classifiers in addition to Ridge classifier and Naive Bayes on CAN-OITDS dataset, where XGBoost achieved the best performance.
+On the other side, the authors of [21] studied ML for V2X anomaly detection using models such as Support Vector Machines (SVM), K-Nearest Neighbors (KNN), Naive Bayes, Random Forest, ensemble boosting, and voting ensemble, and were able to achieve good classification results.
+Despite being lightweight and achieving promising results, simple ML-techniques lack the ability to process the time facet of network data.
+This drawback was addressed by LSTM-based architectures that were designed to effectively model temporal data.
+In  [22, 23, 24], the authors combined LSTM and CNN to derive spatial and temporal features, achieving good misbehavior detection.
+Additionally, the time aspect was further emphasized in [25] where the authors approached misbehavior detection as a time-series anomaly detection problem, modeled as a Markov Decision Process (MDP), where each decision depends on preceding states.
+The improvement introduced by these approaches highlights the importance of the temporal dimension in detecting misbehavior within vehicular networks.
+The emergence of transformers shifted the interest due to their unprecedented capabilities to efficiently process sequenced data types.
+In vehicular networks, CAN-bus anomaly detection took advantage of these capabilities of transformer-based models and showed that BERT-inspired architectures surpass other deep models [18, 26, 27].Beyond works focusing exclusively on CAN-bus, the authors in [28] propose a hybrid Network IDS (NIDS), named IoV-BERT-IDS, for both in-vehicle (CAN-bus) and extra-vehicle networks.
+The model leverages a semantic extractor to convert raw CAN-Bus and network sequences into BERT-compatible representations, pretrained using Masked Byte Word Model (MBWM and Next Byte Sentence Prediction (NBSP) tasks to capture contextual and sequential features.
+It is then fine-tuned for intrusion detection, achieving about 99% accuracy across CICIDS, BOT-IOT, Car-Hacking, and IVN-IDS datasets.
+However, these datasets are not specifically V2X communication datasets; as a consequence, we cannot assume systematically that the model will give the same performance in our study case.
+On the other hand, the authors in [17] focused on V2X communications misbehavior by fine-tuning an encoder-based vanilla transformer in a semi-supervised approach to detect anomalies.
+Their model reconstructs vehicle behavior by regenerating position and speed from sequences of 200 BSMs, then detects misbehavior based on the Mean Squared Error (MSE) between the original and reconstructed features compared to a threshold, yet their data preparation includes a hint to the vehicle’s ID in the dataset, which can lead to data label leakage.
+Additionally, Larger LLM models such as GPT-4o, LLaMA-3.2 and Gemini 1.5 Pro were used in [29] alongside DNNs such as VGG19 and ResNeXt to detect motion forging and natural denoising diffusion (NDD) attacks.
+The results indicate that GPT-4o delivers the strongest performance overall, though the model is not open source, which in a practical scenario will require sharing sensitive vehicular network telemetry with a third party, raising serious privacy and security concerns.
+Furthermore, LLMs were coupled with explainable AI technics explainable AI technics such Kappa coefficient, Matthews Correlation Coefficient, and SHapley Additive exPlanation (SHAP), in [30], to validate the model’s misbehavior detection capabilities.
+This approach studies aims to understand the reason behind a vehicle behavior classification achieved by LLM, however the SHAP results strongly suggest that the message ID feature has the biggest influence on the classification, raising potential overfitting or train/test data leakage.
+From another perspective, the authors in [31] propose a transformer-based Intrusion Detection System (IDS) for V2X trained within federated learning approach.
+Their model is a tailored encoder-based transformer that integrates feature attention and routing to select the most suitable feature set for network packet classification.
+The proposed system overlooks temporal context of the packet in classification, which can be considered a major drawback.
+This limitation is addressed in [32], where the authors train their transformer encoder-based model to classify BSM-windows, showing that the federated learning gives a model with better classification accuracy compared to the standalone model.
+Despite the growing use of LLMs for misbehavior detection, many studies provide limited details on data preparation.
+The results of some works strongly suggest potential training data leakage, particularly when identifiers such as message ID or pseudo-ID features are used, making it difficult to attribute performance gains solely to the model’s learning capacity.
+Moreover, it remains crucial to investigate the computational resources required effectively to deploy such models in real-world settings.
 III Background
 III-A Vehicular Communication Models
-V2X communications encompass a variety of technologies facilitating real-time information exchange between vehicles and their surroundings, thereby enhancing road safety, and traffic efficiency.
-Figure 1 provides an overview of the different types.
+V2X communications encompass various technologies facilitating real-time information exchange between vehicles and their surroundings, thereby enhancing road safety, and traffic efficiency.
+Fig.1 provides an overview of the different types.
 V2V communication involves vehicles equipped with network modules, called onboard units, enabling them to exchange data concerning their position, speed, heading, and acceleration.
 This type of communication aims to prevent accidents by facilitating situational awareness among vehicles.
-In contrast, V2I communication allows vehicles to interact with roadside units (RSUs) such as streetlamps, RFID readers, lane markers, and signage.
+In contrast, V2I communication allows vehicles to interact with RSUs such as streetlamps, RFID readers, lane markers, and signage.
 This bidirectional communication enables the real-time exchange of information about road conditions, traffic, accidents, and more, thereby enhancing overall traffic management and safety[33].
-V2P, or vehicle-to-pedestrian communication, involves interactions with vulnerable entities in the environment, such as pedestrians and cyclists.
+V2P, or vehicle-to-pedestrian communication, involves interactions with Vulnerable Road Users (VRU), such as pedestrians, cyclists and motor wheelers [34].
 These individuals typically carry smartphones or wearable devices capable of transmitting their location to alert nearby vehicles of their presence, contributing to enhanced safety for all road users.
 Additionally, V2X includes V2N communication, which enables vehicles to connect to cloud-based services and platforms.
-This connectivity facilitates access to a wide range of services such as over-the-air updates, remote diagnostics, real-time traffic and navigation information, and data analytics for traffic management purposes.
+This connectivity facilitates access to a wide range of services, such as over-the-air updates, remote diagnostics, real-time traffic and navigation information, and data analytics for traffic management purposes.
 In order to protect ensure vehicle privacy, vehicles use pseudonyms for communications.
 Vehicle pseudonyms are temporary, randomly generated identifiers used in vehicular communication systems to provide privacy and anonymity for vehicles while still allowing them to communicate with each other and infrastructure.
-These pseudonyms are periodically changed to prevent tracking of vehicles over time.
+These pseudonyms are periodically changed to prevent tracking of vehicles over time [35].
 Figure 1: V2X communication and Misbehavior detection with edge and cloud LLMs
@@
 The transformer is a groundbreaking architecture that revolutionized the task of sequenced data processing, introducing innovative mechanisms for capturing complex dependencies and patterns within sequences.
-This was achieved by leveraging the self-attention mechanism to weight the importance of each token in sequence with regard to the other tokens.
+This was achieved by leveraging the self-attention mechanism to weight the importance of each token in sequence compared to the other tokens.
 The weights are obtained by applying three linear transformations, namely (keys (k), queries (Q) and values (V)) on the input sequence of size $d_{k}$ following the equation 1.
+$\text{Attention (Q, K,V)}=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
 Several variations of transformers have been proposed and customized to address a broad spectrum of tasks, encompassing language understanding, generation, translation, and more.
-Notable examples include the original Transformer architecture [30] which is presented in Figure 2, comprised of both encoder and decoder blocks, as well as BERT [12] and RoBERTa [31], both of which are encoder-only transformers.
-Additionally, LLAMA [11] and GPT exemplify the decoder-only type of transformers, primarily designed for generation tasks.
-Mistral-7B, is also a decoder-only transformers that employs specific design decisions, allowing it to outperform models such as the 13 billion-parameter variation of Llama-2 and the 34 billion-parameter Llama-1 with only 7 billion parameters.
-Those technical choices involve using Sliding Window Attention (SWA) that operates by defining a fixed-size window surrounding each token, thus limiting the number of tokens it has to tend for [32].
+Notable examples include the original Transformer architecture [36] which is presented in Fig.
+2, formed by both encoder and decoder blocks.
+BERT [13] and RoBERTa [37], are another type of architecture formed only by encoder blocks.
+On the other hand, LLAMA [12] and GPT are decoder-only types of transformers, primarily designed for generation tasks.
+The difference between the encoder-based and decoder-based transformer is self-attention module type.
+While in the case of BERT-like models, the attention is bidirectional, meaning that all the sequence tokens are used to compute the attention score for each token.
+The decoder-based models use masked self-attention technique, in which computing the attention score of a token requires only past tokens
+Mistral-7B is another example of a decoder-based transformer that leverages specific design decisions, allowing it to outperform models such as the 13 billion-parameter variation of Llama-2 and the 34 billion-parameter Llama-1 with only 7 billion parameters.
+Those technical choices involve using Sliding Window Attention (SWA) that operates by defining a fixed-size window surrounding each token, thus limiting the number of tokens it has to tend to [38].
 Compared to the vanilla self-attention used in [36], this technique enables longer sequence processing with lower computation cost.
-Moreover, Mistral-7B employs the Grouped-query attention (GQA) technique where a group of queries shares a single key and value heads[33], thus speeding up the inference time of the model.
+Moreover, it employs the Grouped-query attention (GQA) technique, where a group of queries shares a single key and value head [39], thus speeding up the model inference time.
+Mistral-7B is suited for misbehavior detection in vehicular networks because it combines the performance of larger models with a more compact size.
+Its smaller size and efficient sequence processing make it more suitable for deployment at the edge, where computational resources are often limited [40].
+Additionally, Mistral-7B’s pretraining on a large and diverse dataset enables easier fine-tuning for specific tasks like ours, requiring less data for adaptation.
+This advantage reduces the overall cost of supervised learning while leveraging its powerful ability to extract semantic information from sequences.
+Overall, Mistral-7B’s capabilities make it suitable for misbehavior detection in vehicular networks.
+Figure 2: Transformer architecture
+III-C Low-rank Adaptation
+In traditional fine-tuning, the model’s weights are directly modified to adapt the model to the new considered task.
+While fine-tuning is a crucial approach in ensuring transfer of knowledge, the process becomes computationally daunting as the number of parameters increases.
+Notably, fine-tuning large language models poses significant challenges in terms of memory requirements and computational resources.
 […]
-Figure 2: Transformer architecture
-III-C Low rank adaptation
-In traditional fine-tuning, the model’s weights are directly modified to adapt the model to the new considered task.
-While fine-tuning is a crucial approach ensuring transfer learning, the process become computationally daunting as the number of parameter increases.
-Notably, fine-tuning large language models poses significant challenges in terms of memory requirements and computational resources.
-Low Rank Adaptation (LoRA), is a technique that was proposed in [34] to alleviate the memory footprint and speed up the process of large models’ fine-tuning.
-This is achieved through reducing the number of trainable parameters by decomposing the changes on the pre-trained weight matrices into low rank matrices.
-The process suppose that after fine-tuning the new weights matrices will have the form $W=W+\Delta W$ where W is frozen during training and $\Delta W$ is decomposed into two lower rank matrices A and B.
-Being of lower rank, the number of trainable parameters in these matrices is smaller than the original $\Delta W$.
 Although LoRA enhances significantly the computational requirements, further optimization could be achieved by using QloRA [42], which applies quantization techniques to further reduce the memory footprint.
@@
 Our method involves reframing the misbehavior detection problem as a sequence classification task.
-In this framework, we deploy a model based on Mistral-7B, which is quantized and implemented on a roadside unit (RSU).
+In this framework, we deploy a model based on Mistral-7B, which is quantized and implemented on a RSU.
 Through quantization, the memory requirements for running Mistral-7B are significantly reduced, enabling its deployment on edge equipment at the RSUs.
 These RSUs possess the necessary capabilities to collect BSMs exchanged within their sensing ranges and execute the model to identify potential misbehavior.
-The collected messages undergo the prepossessing steps presented in Figure 3-a, where the irrelevant features are deleted, the messages are grouped by sender pseudonym and then transformed into BSM-Windows of size $n$ for each user.
-Following the prepossessing steps, the formed time windows are transformed into textual prompts necessary for the language model inference.
-The prompt design steps transform the multivariate time window to a text, as described in Figure 3-b.
-The obtained prompt is then fed to a Tokenizer and then Mistral-7B-based classifier for the final labeling.
-The last step following the behavior assessment is the reporting phase, where the RSU communicates the information’s on the misbehaving vehicle to a trust authority (TA) located in the cloud.
-The TA is then responsible for verifying the report, and feeding it, if necessary, to the cloud LLM, which can complement the report by conducting deeper analysis and validation, and incorporating additional information from a broader dataset stored in the cloud.
+The process of classifying vehicle behaviors consists of three steps, as shown in Fig.3.
+The first step involves data preprocessing, detailed in submodule (a) of Fig.3.
+In this stage, features are extracted from detected BSMs in the RSU’s local database, following the data preparation steps detailed in V-A.
+Thereafter, the BSM messages are grouped by sender pseudonym and transformed into BSM-Windows of size $n$ for each vehicle.
+The second step, represented by the prompt design module in Fig.3, involves converting multivariate time windows into a textual format.
+The text structure is as follows:
+reported values: $Feature_{1}$:[value_1,value_2,,value_n][value\_1,value\_2,...,value\_n], $Feature2$: [value_1,value_2,,value_n][value\_1,value\_2,...,value\_n], ..., $Feature_{M}$: [value_1,value_2,,value_n][value\_1,value\_2,...,value\_n]
+The third and final step is classification, handled by the MistralBSM model in submodule (c) of Fig.3.
+This module consists of a Tokenizer, Mistral-7B layers for prompt processing, and a classification head that assigns a behavior label.
+Following the classification, the RSU reports the information about misbehaving vehicles is communicated to a Trust Authority (TA) located in the cloud.
+The TA is, thus, responsible for verifying the reported misbehavior.
+If further analysis is needed, the TA forwards the report to a cloud-based LLM, which enhances the report by conducting deeper analysis and incorporating additional data from a broader dataset stored in the cloud.
+For instance, the cloud LLM can access historical data on the vehicle’s behavior, as well as sensor data from roadside infrastructure such as LIDAR or radar, allowing for a more thorough assessment of whether the vehicle was transmitting accurate information.
+This deeper analysis is implemented at the cloud level rather than at the edge for privacy reasons, since only the TA has access to the vehicle’s real identity.
+By performing this task in the cloud, sensitive information about the driver or vehicle remains protected and is only accessed when necessary, such as in cases where misbehavior needs to be verified before revoking network access.
 Subsequently, the TA takes the appropriate response that could include revoking the vehicle’s access to the network.
-It is important to highlight that our methodology ensures the examination of both contextual and content aspects within the messages at the edge LLM.
+In this paper, we assume the existence of such a cloud-based model and focus our study on the edge-LLM operating at the RSU.
+In fact, MistralBSM’s methodology ensures the examination of both contextual and content aspects within the messages at the edge LLM.
 Contextual dimensions are considered during the grouping of messages into time windows, facilitating the analysis of temporal patterns and relationships.
@@
 We start by presenting the preprocessing steps we applied on the database used, namely the VeReMi dataset.
-Thereafter, we define the performances metrics used to evaluate the performance of our model.
-After this, we present the two major sets of experiments, binary and multiclass classification.
+Thereafter, we define the performance metrics used to evaluate our model.
+Thereafter, we present the two major sets of experiments: binary and multiclass classification.
 For each set, we investigate two possible methods of leveraging Mistral-7B: one involving fine-tuning and the other without.
-Our aim is to assess whether we can leverage Mistral-7B with minimal modifications, maximizing its utility without extensive adjustments.
+We aim to investigate the possibility of leveraging Mistral-7B with minimal modifications, maximizing its utility without extensive adjustments.
 From there, we compare its performance against two other LLMs, LLAMA2-7B and RoBERTa.
-Lastly, we analyze the effect of varying window sizes on the inference time to identify the necessary requirements and potential solutions for deploying the models.
+Lastly, we analyze the effect of varying window sizes on the inference time to identify the requirements and potential solutions for deploying the models.
 V-A Extended VeReMi Dataset
@@
 It was generated using the Framework For Misbehavior Detection (F2MD), which is an extension of the vehicular network simulator VEINS, and a subsection of the Luxembourg SUMO Traffic (LuST) scenario covering 1.61 km² with a peak density of 67.4 vehicles per km².
-To experiment and assess our approach, we used the extended version of VeReMi[37] that offers more diverse attack types including DoS, Random Positioning Attacks, Constant Positioning Attacks, Disruptive Attacks, Constant Speed Attacks, Delayed Messages Attack, Data Replay Attacks, Traffic Sybil Attacks, and Eventual Stop Attacks that were simulated under varying traffic conditions specifically from 07 to 09 am and from 2 to 4 pm.
+To experiment and assess our approach, we used the extended version of VeReMi[44] that offers more diverse attack types, including DoS, Random Positioning Attacks, Constant Positioning Attacks, Disruptive Attacks, Constant Speed Attacks, Delayed Messages Attacks, Data Replay Attacks, Traffic Sybil Attacks, and Eventual Stop Attacks that were simulated under varying traffic conditions, specifically from 07 to 09 am and from 2 to 4 pm.
 The dataset is structured into trace files containing the exchange of BSMs by individual vehicles.
@@
 Duplicate entries resulting from the interception of the same message by multiple vehicles were eliminated.
-Then fields describing noise attributes such as position noise, heading noise, acceleration noise and speed noise were dropped.
-In order to minimize memory usage during fine-tuning and inference, adjustments were made to the structure of certain fields.
-The heading, position, acceleration, and speed fields, originally consisting of lists of three values each representing the x, y, and z coordinates, underwent transformation.
-For speed and acceleration, the norm was calculated, condensing the data into single values.
+Then fields describing noise attributes such as position noise, heading noise, acceleration noise, and speed noise were dropped.
+In order to optimize memory usage during fine-tuning and inference, we made adjustments to some fields.
+For instance, we calculated the Euclidean norm for speed and acceleration fields in order to compress the 3-value features into a single value.
 Additionally, the position and heading fields were split into two distinct fields: position-x and position-y, and heading-x and heading-y, respectively.
@@
 Lastly, to further reduce memory requirements, all numerical values were rounded to three decimal points.
-Following this, each message was labeled based on its sender ID, therefore, every message issued by a misbehaving car was labeled as attack.
+Thereafter, each message is labeled based on its sender ID; therefore, every message issued by a misbehaving car was labeled as attack.
 This classification is derived from the title of the corresponding trace file, referencing the attack-code correspondence outlined in Table I.
-Once the dataset was labeled, messages for each sender ID were grouped, ordered by the receive time and segmented into time windows.
-The final preprocessing step, is the textualization of the time windows, which could be achieved either by considering each row as a separate sentence having the feature values preceded by the corresponding field name, or column wise like described in Fig.
-3-b.
+Once the dataset was labeled, messages for each sender ID were grouped, ordered by the receive time, and segmented into time windows.
+Importantly, we removed the sender ID, sender Pseudo, message ID features to avoid label leaking during training.
+The final preprocessing step is the textualization of the time windows, which could be achieved either by considering each row as a separate sentence having the feature values preceded by the corresponding field name or column-wise, like described in the submodule (b) in Fig.
+3.
 As the former results in large prompt size due to repeated field names, we resorted to the column-wise approach.
-For our experiments, We selected a subset of attacks from the VeReMi dataset, and we processed them to extract time windows.
-In order to be able to select what is the best number of packets inside a time window, we created different datasets with 4 sizes, 10, 20, 50 and 100.
-Table I and Figure 4 give an overview of the selected attacks and class distribution after processing.
+For our experiments, we consider the two prominent types of misbehaviors in V2X communications: faulty-hardware-like behavior and purely malicious attacks.
+Therefore, we select Constant Position, Constant Speed, and Eventual Stop as the former type and Data Replay, Dos, Dos Random, and Dos Random Sybil as the malicious attacks from the extended VeRiMi dataset, and we preprocess them to extract BSM sequences as time windows.
+TABLE I: Mapping attack type to codes
 Identifier Attack Type Attack description
@@
 A18 DoSRandomSybil Malicious node would use Sybil attack to conduct DoS random.
-TABLE I: Mapping attack type to codes[37]
+In order to be able to select what is the best number of packets inside a time window, we created different datasets with 4 sizes, 10, 20, 50 and 100.
+Table I and Fig.4 give an overview of the selected attacks and class distribution after processing.
 Figure 4: The class distribution after dataset preprocessing
@@
 A high recall reflects the detector’s ability to identify the actual attack and minimize the chances of labeling an attack as normal.
-Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ (2)
+$\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ (2)
 Precision
 Calculates the correctly labeled attacks to the actual attacks, forming an idea on the model’s capacity to identify attacks while minimizing false alarms.
-Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$ (3)
+$\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$ (3)
 F1 score
@@
 Computes the percentage of instances correctly classified by the model. It reflects the model’s ability to correctly identify the normal from the anomalous behavior.
-Accuracy=True Positives+True NegativesTotal $\text{Accuracy}=\frac{\text{True Positives}+\text{True Negatives}}{\text{TotalInstances}}$ (5)
-V-C Binary classification
+$\text{Accuracy}=\frac{\text{True Positives}+\text{True Negatives}}{\text{Total Instances}}$ (5)
+V-C Misbehavior Detection
 In the first set of experiments, we utilize Mistal-7B to distinguish between normal behaving cars and misbehaving ones.
@@
 Furthermore, the $r$ and the LoRA alpha in the Table II refer to the dimension of the low-rank matrices and the scaling factor which calibrates the weights update magnitude during the training, respectively.
-Figure 5 presents the effect of the hyperparameter r on the accuracy when window size equals 10.
+Fig.
+5 presents the effect of the hyperparameter r on the accuracy when window size equals 10.
 The figure shows that the best accuracy is obtained for an r value of to 2, whereas it decreases significantly for higher values.
@@
 Notably, this was achieved despite the utilization of only 1000 instances per class and tuning only 0.012% of the parameters.
+TABLE II: Hyperparameter settings for fine-tuning with QLoRA
 Hyperparameter Values
@@
 Task type sequence classification
-TABLE II: Hyperparameter settings for fine-tuning with QLoRA
+TABLE III: Binary Mistral-7B results
 Window size Accuracy F1 Recall Precision
@@
 100 0.980 0.98020 0.980120 0.98028
-TABLE III: Binary Mistral-7B results
 V-C2 Mistral-7B as Feature Extractor
@@
 Since we are taking the maximum over the tokens’ representation on each model dimension, the larger the sequence, the higher the chance of loosing information.
+TABLE IV: Binary classification head hyperparameters
 Hyperparameter Value
@@
 Batch size 32
-TABLE IV: Binary classification head hyperparameters
+TABLE V: Binary MistralFeatureNet results
 Window size Accuracy F1 Recall Precision
@@
 100 0.820000 0.900000 0.860000 0.85000
-TABLE V: Binary MistralFeatureNet results
-V-C3 Other LLMs for binary classification
+V-C3 Other LLMs for Attack Detection
 To further validate the model choice, we fine-tuned RoBERTa and LLAMA2-7B under similar conditions.
@@
 With LLAMA2-7B we obtained good results with only 1000 instance per class, whereas RoBERTa required a minimum of 3000 per class to converge.
-Figure 6 summarizes the different models performance across the time windows.
+Fig.
+6 summarizes the different models performance across the time windows.
 The immediate observation is the superiority of fine-tuned Mistral-7B over all the other models on most metrics, notably the accuracy and the F1 score.
@@
 Another contributing factor to this behavior is the textualization step described in Section V-A, in which each column is considered as a sentence, so the truncation step results in eliminating whole features from the sequence.
-Finally, Figure 6 suggests that 100 is the most suitable window size for the fine-tuned Mistral-7B.
+Finally, Fig.
+6 suggests that 100 is the most suitable window size for the fine-tuned Mistral-7B.
 In contrast, 50 is the most suitable for LLAMA2-7B and MistralFeatureNet as they exhibit a deterioration when the window size is increased further.
 Figure 6: Exploring Classification Metrics Variability among Binary Models
-V-D Multiclass classification
-For multi-class classification, the goal is to use Mistral-7B to detect and correctly classify attacks among the 9 classes described in Table I.
+V-D Misbehavior Classification
+For misbehavior classification, the goal is to use Mistral-7B to detect and correctly classify attacks among the 9 classes described in Table I.
 We employ the identical reasoning as in the binary categorization.
@@
 In this case, we select 1000 samples of each class to ensure the balance between the classes.
-Figure 8 shows the performance results for different window sizes.
+Fig.
+8 shows the performance results for different window sizes.
 As with binary classification, the hypothesis that the model performs better with a larger window is validated.
 The refined model obtained an F1 score of 0.96 in a 9-class classification task with just 1,000 instances per class.
-The confusion matrix in Figure 7 also proves the potential of the model, which correctly discerns most classes.
+The confusion matrix in Fig.
+7 also proves the potential of the model, which correctly discerns most classes.
 Figure 7: Mistral-7B Confusion Metrics
@@
 This is why we retained the fine-tuned.
+TABLE VI: Classification network hyperparameters
 Hyperparamter Value
@@
 Batch size 64
-TABLE VI: Classification network hyperparameters
+TABLE VII: Multi-class MistralFeatureNet results
 Window Accuracy F1 Score Recall Precision
@@
 100 0.6128 0.5654 0.6128 0.5587
-TABLE VII: Multi-class MistralFeatureNet results
-V-D3 Other LLMs for multiclass classification
-The performance of LLAMA2-7B and roBERTa on the nine-class classification task is illustrated in Figure 8.
+V-D3 Other LLMs for Attack Classification
+The performance of LLAMA2-7B and RoBERTa on the nine-class classification task is illustrated in Fig.
+8.
 During training, LLAMA2-7B was configured similarly to Mistral-7B, with each class containing 1000 instances.
@@
 V-E Inference Time and Deployment Considerations
-Due to the size of the Mistral-7B model, deploying it requires a balance between multiple hyperparameters to avoid adding overhead to a process that requires speed and accuracy.
-One method to mitigate memory demands is through quantization.
-In our experiments we used 4-bit quantization, which reduces significantly the memory required to load the model to 4GB.
-In addition to memory requirements, it is also very important to study the inference latency.
-It is evident that the size of the prompt is one of the main factors affecting inference time.
-While we demonstrate that the best window size for binary and multiclass classification is 100 using fine-tuned Mistral-7B, opting for these values might introduce inference latency which is limitating to the global performance of the MDS.
-Figure 9, illustrates the evolution of the inference time with regard to the window size for Mistral-7B and LLAMA2-7B.
-We observe similar trends for both models, where the inference time increases notably as the window-size increases.
-This patern may pose a constraint on the use of LLMs in this context.
-Importantly, these measurements were conducted on A100 40GB GPUs and depending on the hardware, actual values may vary from the reported figures.
-However, these models are, designed to seamlessly process prompts with different sizes without having to modify the architecture or apply additional preprocessing, which guarantees flexibility in terms of the window-size.
-Therefore, based on the computational resources available, the size can be adapted to enhance performance without notable accuracy reduction.
-the figures have shown that Mistral-7B.
-maintains good performance even with smaller window sizes.
+[…]
+On the other hand, inference time is influenced by many factors, including the type of computational hardware, the size of the model, and the prompt length.
+In our case, the prompt is the BSM time window.
+Therefore, we focused our analysis on the impact of the window size on detection performance.
+[…]
+Despite the latency limitation, MistralBSM is designed to handle variable prompt lengths without requiring architectural modifications or additional preprocessing, ensuring flexibility in adapting the window size.
+As a consequence, the window size can be adjusted based on available computational resources to optimize performance without significant loss in accuracy.
 Figure 9: Inference Time per Window Size